Karlo is now four times as old as Kevin. In 12 years, Karlo will be twice as old as Kevin. How old are Karlo and Kevin?
Answer by lwsshak3(11628)

Karlo is now four times as old as Kevin. In 12 years, Karlo will be twice as old as Kevin. How old are Karlo and Kevin?
**
let x=Kevin's present age
4x=Karlo's present age
12 yrs hence:
(x+12)=Kevin's age
(4x+12)=Karlo's age
4x+12=2(x+12)
4x+12=2x+24
2x=12
x=6
4x=24
Kevin's present age=6
Karlo's present age=24